Solve the equation for y 2x+10y=3 Trying to find the slope by solving the equation for y
Answer by Alan3354(69443)

Solve the equation for y 2x+10y=3 Trying to find the slope by solving the equation for y
--------------------
2x + 10y = 3
10y = -2x + 3
y = (-1/5)x + 3/10
slope, m = -1/5